The sum of the digits of a two-digit number is 6. When the digits are reversed, the resulting number is 6 greater than 3 times the original number. Find the original number.
Answer by stanbon(75887)

The sum of the digits of a two-digit number is 6. When the digits are reversed, the resulting number is 6 greater than 3 times the original number. Find the original number.
----
Original Number: 10t+u
---
Equations:
t + u = 6
10u+t = 3(10t+u)+6
-----
Simplify:
10u+t = 30t+3u+6
7u-29t = 6
-------
Substitute for "u" and solve for "t":
7(6-t)-29t = 6
42-36t = 6
36t = 36
t = 1
---
u = 6-t = 5
--
Original Number: 15
==========================
Cheers,
Stan H.
==========================